OpenStudy (anonymous):

Explain, how you would use the substitution method to solve the following system of equations. x + 6y = 1 2x + 5y = 9

OpenStudy (anonymous):

x=1-6y so 2x+5y=2-12y+5y=2-7y=9 ==> y=7/-7=-1 Thus x = 1+6 = 7 If that's what you call substitution method